If Shantel walks 4 miles in 50 minutes, then Shantel will walk how far in 90 minutes if she walks at the same speed the whole time? If necessary, round your answer to the nearest tenth of a mile.

Answer by stanbon(75887)

If Shantel walks 4 miles in 50 minutes, then Shantel will walk how far in 90 minutes if she walks at the same speed the whole time? If necessary, round your answer to the nearest tenth of a mile.
-----
Use a proportion::
w/90 = 4/50
w = (9/5)4
w = 36/5 = 7.2 miles
---------------
Cheers,
Stan H.
-------------